A rectangular swimming pool is 11.5 m long and 5 m wide. How many metres of fencing are needed to enclose the pool if you erect the fencing 2m from the edge? Show your calculations:
Answer by ikleyn(52781)

.
A rectangular swimming pool is 11.5 m long and 5 m wide.
How many metres of fencing are needed to enclose the pool
if you erect the fencing 2m from the edge? Show your calculations
~~~~~~~~~~~~~~~~~

The length of the fence is 2+2 = 4 meters longer than the length of the pool,

so it is 11.5 + 4 = 15.5 meters.



The width of the fence is 2+2 = 4 meters longer than the width of the pool,

so it is 5 + 4 = 9 meters.



Now compute the perimeter of the rectangular fence  15.5 + 9 + 15.5 + 9 = 49 meters.


ANSWER.  49 meters of fencing is needed.

Solved.